Circular No. 8590

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 Q1 (LINEAR) An apparently asteroidal object discovered by the LINEAR project

(discovery observation given below), and posted on the ‘NEO Confirmation Page’, has been found cometary on CCD images taken by several observers. J. Lacruz (Madrid, Spain, 0.30-m reflector) found a fan-shaped tail 12′′ long in p.a. 200^{o} on his co-added images from July 28.03–28.05 UT. J. E. McGaha (Tucson, AZ, 0.62-m reflector) found a soft 9′′ coma with a slight elongation to the southwest on his images from Aug. 28.4. J. Young (Table Mountain, 0.6-m reflector) found a 6′′ coma with no central condensation on his images taken on Aug. 28.4 and 29.5, with a 16′′-long fan-shaped tail spanning p.a. 190^{o}-230^{o} on the second date (moonlight, poor seing, and nearby stars preventing a view of the tail on the first night).

2005 UT  α_{2000}  δ_{2000}  Mag.
Aug. 27.41707  5^{h}23^{m}02.^{s}36  +64^{o}38′01′′.3  18.9

The available astrometry, the following uncertain preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2005-Q48.

T = 2006 Mar. 19.437 TT  ω = 63.^{o}404
Ω = 88.927 }2000.0
q = 6.18426 AU  i = 106.456

SUPERNOVA 2005do IN UGC 2495 Further to IAUC 8589, E. Lee and W. Li report the LOSS discovery

of an apparent supernova on unfiltered KAIT images taken on Aug. 8.50 (at mag 17.4), 11.47 (mag 17.6), 22.51 (mag 18.1), and 23.46 UT (mag 18.2). SN 2005do is located at α = 3^{h}02^{m}00^{s}.42, δ = +41^{o}34′34′′.2 (equinox 2000.0), which is 70′′.4 west and 62′′.0 south of the nucleus of UGC 2495. A KAIT image taken on 2004 Jan. 30.16 showed nothing at this position (limiting mag ∼ 19.0).

F. J. D. Serduke, A. Blum, J. Scala, and A. V. Filippenko, University of California, Berkeley, report that inspection of a CCD spectrum (range 330– 1000 nm), obtained on Aug. 26 UT with the Shane 3-m telescope at Lick Observatory, shows that 2005do is a supernova of type Ia, ∼ 1–2 months past maximum brightness.

2005 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green